OpenStudy (anonymous):

solve the DE y'''+2y'' +y=0

OpenStudy (anonymous):

let y=e^(mx) be the trial solution now here y' = my and y'' =m^2 y thus the auxillary eq is m^2+m+1=0 or (m+1)^2 =0 because roots of above eq is -1,-1 thus the required solution is y=(a+bx)*e^(-x)

OpenStudy (anonymous):

wouldn't the auxillary eq be m^3 +m^2 +1=0?

OpenStudy (anonymous):

yup u r correct i mistook the question.